Kropf Inc. has provided the following data concerning one of the products in its standard cost system. Variable manufacturingMaterials price variance a. b. Materials quantity variance Labor rate variance C. d. Labor efficiency variance eVariable overKropf Inc. has provided the following data concerning one of the products in its standard cost system. Variable manufacturing overhead is applied to products on the basis of direct labor-hours. Inputs Standard Quantity or Hours per Unit of Output Standard Price or Rate Direct materials 8.60 liters $ 8.20 per liter Direct labor 0.70 hours $ 26.70 per hour Variable manufacturing overhead 0.70 hours $ 7.10 per hour The company has reported the following actual results for the product for September: Actual output 10,800 units Raw materials purchased 94,200 liters Actual cost of raw materials purchased $ 782,500 Raw materials used in production 92,900 liters Actual direct labor-hours 7,110 hours Actual direct labor cost $ 193,302 Actual variable overhead cost $ 45,414 Required: a. Compute the materials price variance for September. b. Compute the materials quantity variance for September. c. Compute the labor rate variance for September. d. Compute the labor efficiency variance for September. e. Compute the variable overhead rate variance for September. f. Compute the variable overhead efficiency variance for September. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1

1) Material price variance = (8.2*94200-782500) = 10060 U

2) Material quantity variance = (10800*8.6-92900)*8.2 = 164 U

3) Labor rate variance = (26.70*7110-193302) = 3465 U

4) Labor efficiency variance = (10800*.7-7110)*26.70 = 12015 F

5) Variable overhead rate variance = (7.1*7110-45414) = 5067 F

6) Variable overhead efficiency variance = (10800*.7-7110)*7.1 = 3195 F